### v2

# secReferences in the Online Appendix

Abstract
We examine the growing gender gap in venture capital funding, focusing on accelerator programs in the U.S. We collect a unique dataset with detailed information on accelerators and startups. Using a two-stage methodology, we first estimate a matching model between startups and accelerators, and then use its output to analyze the gender gap in post-graduation outcomes through a control function approach. Our results suggest that female-founded startups face a significant funding disadvantage due to relocation challenges tied to family obligations. However, larger cohorts and higher-quality accelerators help reduce this gap by potentially offering female founders better networking opportunities and mentorship.
Keywords: venture capital, business accelerator, gender, startup, structural matching model
JEL Codes: L1, L26, J16, G24
1 Introduction
Over the past decade, the share of venture capital (VC) deals involving women-founded startups in the U.S. has steadily increased. However, the funding gap has widened in dollar terms, with women-founded startups receiving smaller average investments. Panel (a) of Figure 1 illustrates that while the gender gap in deal frequency has narrowed (blue dashed line), the disparity in funding amounts has grown (red solid line). This funding gap appears to be driven by late-stage investments (gray dashed line), which tend to be larger in size than earlier-stage funding (Panel b). In this paper, we explore the trend in the funding gap, discuss its underlying causes, and examine potential solutions.
Figure 1: The gender gap in venture capital funding
(a) The gap in US\$ is widening
(b) The gap rises in early and late stage
Notes: Using data from Pitchbook, (2021, 2022), we compute VC deal counts and average deal sizes by gender from 2011 to 2020. Panel (a) shows two measures of the gender gap in VC funding: the red line represents the dollar-valued gap in deal size between male- and female-founded startups, while the blue dotted line plots the difference in deal frequency between all-male and women-founded startups (left vertical axis). Panel (b) plots the relative funding gap across VC stages, computed as $1-\frac{\text{avg. funding for women-founded}}{\text{avg. funding for all-men founded}}$. Lines are regression fits by group. A startup is “women-founded” if at least one founder is a woman but results are qualitatively similar with other classifications.
Collecting detailed data on startups is challenging. Startups are more difficult to track compared to publicly listed firms, and, more importantly, there is often limited information on their operations. Who are the founders? What are their gender and background? How much funding have they raised? These challenges complicate productivity comparisons between startups led by female and male founders, as it becomes nearly impossible to control for the quality of a startup, which is crucial to identify discrimination.
We address these problems by studying startups graduating from accelerators. Accelerators can be seen as “colleges for startups,” as they provide services (e.g., legal support) and mentorship. Similar to college admissions, startups apply to multiple accelerators before enrolling in one of them. We hand-collected information on the universe of startups “graduating” from all U.S. accelerators from their respective websites and other sources such as Linkedin. For each startup, we observe demographic information regarding its founders, funding outcomes over different horizons since graduation, as well as whether they were acquired or went public. We observe various characteristics for accelerators as well, such as their yearly cohort.
Because accelerators selectively admit startups based on potential and because startups apply to accelerators based on fit, the match between a startup and its accelerator contains valuable information about the quality of each enrollment match. In fact, standard regression approaches that ignore this selection process risk conflating founder gender with latent quality differences. To address this, we adopt a two-step empirical strategy that explicitly models the startup–accelerator matching process (e.g., Sørensen, , 2007).
In the first step, we estimate a one-to-many matching model with non-transferable utility between startups and accelerators. This model captures the assortative nature of admissions: high-quality startups tend to match with high-quality accelerators. Importantly, our framework allows startups to match with accelerators outside their home region, enabling us to analyze geographic mobility. The matching process is competitive and capacity-constrained, with equilibrium determined by pairwise stability (Roth and Sotomayor, , 1990). [^2] This approach differs substantially from the standard discrete choice approach. In a discrete choice model where startups select accelerators, accelerators do not face capacity constraints. As a result, the market competition encountered by startups is not accounted for, which can lead to biased estimates. The resulting matching probabilities provide a proxy for unobserved startup quality and the competitive intensity of the admissions process.
In the second step, we analyze post-graduation VC funding outcomes using a control-function approach. By incorporating the estimated match-level information from the first step, we control for endogenous selection into accelerators. This allows us to isolate the effect of founder gender on funding success, net of both observed and unobserved quality differences (Lu and Rui, , 2018; Akkus et al., , 2020).
Our approach introduces a novel two-stage estimation procedure that separates the matching process from the outcome equation. The key advantage of this separation is that it enables a more encompassing definition of markets, obviating the need to segment the U.S. into multiple local markets. This is particularly important in our setting, where 27.17 percent of startups relocate to accelerators outside their home states, rendering localized market definitions impractical. While this comes at the cost of some statistical efficiency compared to joint estimation, it greatly simplifies the likelihood function in the first stage. With such analytical tractability, we do not need computationally intensive simulations, which would require drawing error terms for an astronomical number of possible matches to evaluate the likelihood function. [^3] For instance, a market with 100 startups, 20 accelerators with 5 slots each has more than 10100 possible matches.
The separation of stages also facilitates the application of a control function approach in the second stage, enabling flexible regression-based analysis across various outcome variables. This structure is computationally convenient: by estimating the first stage only once, we can bootstrap the second-stage standard errors without repeatedly solving the matching problem. Although this two-step procedure sacrifices some asymptotic efficiency relative to fully joint estimation methods (e.g., Sørensen, , 2007), it enables scalable inference in the presence of large market sizes.
We find strong evidence of a significant gender penalty. Startups with a female founder have a -3.8 percentage point lower probability of reaching the five-million-dollar funding milestone one year after graduation, with the gap widening to -16 percentage points within five years. Since later VC rounds tend to be larger than earlier ones, this widening probability gap translates into an even larger gap in U.S. dollar terms. This result is particularly concerning because attending an accelerator serves as a signal of quality, meaning the quality heterogeneity among the startups in our sample should be smaller than the heterogeneity underlying Figure 1, which draws from the universe of U.S. startups.
What mechanisms drive this post-accelerator gender gap? Our analysis points to geographic mobility constraints as a key factor. The disadvantage for female founders is most pronounced among those in their late 20s to early 30s, an age range when many face family obligations or childcare responsibilities. These constraints make relocation more difficult, leading some women to enroll in less prominent local accelerators or to forgo the opportunity altogether. In turn, this limits their access to investor networks and follow-on capital.
At the same time, we find that female-founded startups that do relocate, overcoming family-related location frictions, continue to lag male-founded startups in early fundraising, but nearly close the gap within five years after graduation. This convergence is especially strong when women-led ventures participate in accelerators with large cohorts or top-tier reputations, which likely offer richer mentorship and broader investor exposure. These results suggest that accelerator programs can partially mitigate the gender funding gap: by providing peer support and access to professional networks, they help female entrepreneurs overcome disadvantages linked to mobility constraints and narrower informal connections.
Our findings carry important implications. First, we contribute to research on the mechanisms behind gender disparities in entrepreneurship. While prior work has emphasized behavioral differences (Croson and Gneezy, , 2009), statistical discrimination, or gendered pitch perceptions (Kanze et al., , 2018; Exley and Kessler, , 2022), our findings highlight geographic mobility constraints—often tied to family responsibilities—as an underappreciated barrier for female founders. This complements evidence from labor and urban economics on how commuting costs and caregiving roles shape women’s labor market choices and entrepreneurial activity (Bertrand et al., , 2010; Rosenthal and Strange, , 2012; Le Barbanchon et al., , 2021; Jayachandran, , 2021; Zandberg, , 2021). Crucially, we show that even when women overcome these frictions and access top-tier accelerator programs, funding gaps persist —suggesting that structural constraints, rather than pitch style or ambition, are central to the problem. Our analysis also connects to broader work on the “glass ceiling” in both labor markets and science, where similarly qualified women face persistent barriers to advancement (Bertrand et al., , 2018; Galasso and Profeta, , 2024). [^4] This paper also broadly relates to the extensive literature on the gender gap in the labor market (e.g., Biasi and Sarsons, , 2022; Campa et al., , 2011; Dahl et al., , 2021; Ghazala and Ferrer, , 2017; Kawaguchi, , 2007) and in the sciences, as most of the founders in our datasets have STEM backgrounds (Iaria et al., , 2024; Ahn et al., , 2024; Galasso and Profeta, , 2024).
Second, we show that certain features of the entrepreneurial ecosystem can alleviate this disparity. In our data, female-led ventures are just as likely to survive and to be acquired as their male-led counterparts, despite receiving less early-stage funding (e.g., Ewens and Townsend, , 2020; Gornall and Strebulaev, , 2024). This suggests that capital constraints on women-founded startups represent lost growth opportunities rather than rational responses to poorer performance (Huang and Kisgen, , 2013; Nanda and Rhodes-Kropf, , 2013). Closing the gender gap in financing, therefore, is not only a matter of equity but also one of economic efficiency: high-potential ventures led by women may be underfunded relative to their merits, which hampers innovation and limits new firm entry. More broadly, such underfunding may reduce the variety of business ideas brought to market (Pistilli et al., , 2023).
Finally, our study sheds light on how organizational interventions can promote inclusion. Accelerators—especially investor-driven programs like Y Combinator or TechStars—have become a critical gateway to venture funding (Hallen et al., , 2020; Yu, , 2020). Our results suggest that these programs, through their cohort-based mentorship and investor networks, serve as an equalizing force that can partially counteract gender-based network gaps (Clingingsmith et al., , 2022; Cohen et al., , 2019). This opens up new avenues for research and practice: for example, accelerator managers and policy makers might consider cohort policies or support services (such as child-care assistance or remote program options) to encourage greater participation of female founders who face mobility constraints (Balachandra, , 2019; Brush and Elam, , 2024). In sum, by identifying relocation frictions as a barrier and accelerators as a tool to overcome it, our study contributes a deeper understanding of why female entrepreneurs remain underfunded and what can be done to bridge this gap.
The paper is structured as follows. Section 2 explains the dataset and the data collection process. Section 3 describes the role of accelerators and compares our data to the broader venture capital market. Section 4 presents the matching model and its estimation, with the results detailed in Section 5. Section 6 explores relocation as a key driver of the gender gap, discussing the implications of our findings. Finally, Section 7 concludes the paper.
2 Data
We construct a novel dataset covering all U.S. accelerators from 2008 to 2011. Collecting startup data is inherently challenging. We began by identifying accelerators from seed-db.com, a well-known public repository of accelerator programs. However, its lists of accelerator participants can be incomplete, particularly for less prominent programs. To address this, we supplemented our data using Google News and platforms like TechCrunch, retrieving press releases and announcements from the relevant cohort years. To our knowledge, we have covered all participants of these cohorts.
Most accelerators during this period were investor-led programs focused on IT industries, with many well-known accelerators emerging at this time. We exclude non-profit-driven accelerators, including those with community restrictions, government or non-profit funding, or no equity stake. This ensures our dataset consists only of accelerators maximizing expected returns. We also omit startups with missing characteristics. [^5] Most omitted startups had already ceased operations. However, they do not appear systematically different from other failed startups, and their exclusion, representing about 5% of the dataset, is unlikely to impact our results significantly.
We collect program and participant details from CrunchBase, AngelList, CapitalIQ, CBInsights, VentureXpert, and LinkedIn. Data on private firms often suffer from self-reporting bias, as successful startups are more likely to disclose information. To mitigate this, we cross-check firms using news articles and press releases. However, self-reporting bias is minimal in our dataset, as most startups had publicly available information due to the prominence of accelerators.
We define a “program” as a cohort of startups. Some accelerators operate multiple programs across different locations and time periods. In total, we identify 74 programs from 27 accelerators, covering 736 startup graduates (the average cohort size is 17.7 startups). 18 of these programs have startups with at least a female founder. Around 37% of the programs were located in startup hubs (CA, MA, NY), closely reflecting the broader geographic distribution of U.S. accelerators, with approximately 40% concentrated in major tech hubs such as Silicon Valley, Boston-Cambridge, and New York.
Table 1 presents summary statistics for accelerator participants. [^6] Industries are classified using the six-digit 2012 NAICS codes: IT Service (519190), Software (511210), Data Processing and Hosting (518210), Internet and Web (519130), and Others (e.g., Healthcare, Mobile Devices). Overall, gender differences are small, with differences across most variables being not statistically significant at the 10% level. However, women-founded startups, meaning startups with at least one female founder, have fewer founders with Engineering or Science degrees, smaller founding teams, older founders, and are less likely to operate in Software. We control for these differences in our analysis.
Table 1: Startup Profiles
All Women Founded Men Founded Founding Team Startup age (years) 0.762 0.889 0.750 At least one serial founder 0.376 0.286 0.385 Team size 2.264 2.016 2.287 Average age of founders 28.776 30.630 28.603 At least one graduate degree 0.353 0.429 0.346 At least one Ph.D. 0.075 0.111 0.071 At least one engineering/science degree 0.644 0.317 0.675 Industry IT services 0.382 0.317 0.388 Software 0.186 0.222 0.183 Data processing & hosting 0.268 0.317 0.263 Internet & web 0.092 0.032 0.098 Other industries 0.072 0.111 0.068 Observations 736 63 673 Notes: This table presents summary statistics for the startups in our sample. “Women Founded” refers to startups with at least one female founder. “Men Founded” are startups with all-male founding teams. “Serial founder” refers to a founder who had previously created at least one other startup.
3 Background
3.1 What are accelerators?
Accelerators, also known as “seed accelerators” or “startup accelerators,” primarily target early-stage, high-tech startups, especially in IT-related industries. During our data period, most accelerators admitted startups ready to raise VC, typically taking a 5% equity stake.See Appendix A.1 for details on accelerator operations.
According to Cohen and Hochberg, (2014), investor-led accelerators can serve as deal aggregators for venture investors. The unique structure of accelerators helps VCs select startups by combining the funds of many investors and spreading risk across more portfolio firms. In practice, accelerator fund investors often increase their investments in their favorite startups post-accelerator program.
The startups that have graduated from accelerators in our dataset also attract venture investors who do not directly invest in accelerators. In our accelerator data, over 40% of accelerator graduates received VC immediately after graduation. In comparison, only approximately 3% of high-tech startups ever receive VC investments according to the Kauffman Foundation Survey—a panel study of 4,928 businesses founded in 2004 and tracked over their early years of operation through 2011.
Table 2 compares the funding received by accelerator graduates with that of comparable startups in the VC market. Panel A reports the average deal size across years. For each cohort from 2008 to 2011, the table shows the average funding received one year after graduation and compares it to that of similarly aged startups in the “seed” or “angel” stages of the VC market for the corresponding year (i.e., firms in the very first stages of the VC market). We find that accelerated startups secure slightly more funding in their first year post-graduation, suggesting they are of higher quality than the broader VC market.
Panel B examines their average yearly fundings over a longer horizon of five years. Each year, reports the total five-year funding divided by the number of years in operation for either the accelerated startups in our sample, or a sample of Pitchbook startups in the “seed” to “late” stages of the VC market. Also in this longer horizon, accelerated startups generally receive more funding than the average startup in the VC market, showing that their higher quality is maintained over time.
This comparison highlights the importance of controlling for variation in startup quality in our analysis.
Table 2: Accelerated startups receive more funding than other startups in the VC market
Graduation Year Average Yearly Funding for Cohort: 2008 2009 2010 2011 Panel A: In the first year since graduation Accelerator graduates 0.82 1.37 1.36 1.45 Comparable startups (Seed and Angel stages) 1.06 1.00 0.86 0.87 Panel B: In the first five years since graduation Accelerator graduates 1.55 6.70 2.42 2.06 Comparable startups (same age) 2.64 2.05 1.96 1.95 Values in US\$ millions. Notes: This table compares funding received by accelerator graduates with comparable startups in the VC market. Panel A reports the average deal size for accelerator graduates from 2008 to 2011 in their first year post-graduation, alongside startups of similar age in the “seed” or “angel” stages of the VC market for the corresponding year. Panel B presents the average yearly funding over a five-year horizon, calculated as total funding over five years divided by years in operation. The comparison sample consists of startups in the “seed” to “late” stages from PitchBook. Funding amounts are inflation-adjusted to constant dollars.
3.2 Accelerators’ Admission and Gender Diversity
Accelerators were relatively new in the venture market, with approximately 800 graduates by the end of 2011. Gender diversity was not a selection criterion at the time. For instance, Stross, (2012) notes that Y Combinator focused solely on startup growth potential and did not attempt to balance gender in applications. Criticism for admitting too few women only emerged post-2012. Y Combinator, often scrutinized for its low admission rate of women-founded startups, faced such criticism in 2013.777See: http://www.paulgraham.com/ff.html.
To test whether public opinion influenced accelerator admissions over our sample period, we regress the probability of admitting a women‐founded startup on state‐level vote share for the Democratic presidential candidate in 2008, following Giuli and Kostovetsky, (2014). The hypothesis is that accelerators located in states with stronger Democratic support would be more sensitive to gender bias. As shown in Appendix Table A1, we find no statistically significant relationship between Democratic vote share and the likelihood of admitting women‐founded startups, suggesting that there is no evidence that accelerators favored women founders during our sample period.
Therefore, we rule out the possibility that accelerators showed systematically greater favoritism toward certain types of startups—such as those founded by women—in a way that would alter the interpretation and external validity of our results with respect to the gender gap in financing.
3.3 The Gender Gap in Startup Performance
We define a startup as “women-founded” if at least one of its founders is female. While this definition is broad, it is necessary given the scarcity of startups with all-women founding teams. Using data on the universe of U.S. startups from Pitchbook, (2021, 2022), Appendix A.2 shows that similar patterns to the ones we discuss in this paper emerge under alternative definitions of “women-founded,” including those restricted to all-women teams. As a result, our estimates should be interpreted as a lower bound on the true gender gap in startup founding—both due to the inclusive definition of “women-founded” and because accelerators tend to attract the most promising startups.
Table 3 compares the performance post-accelerator of women- and men-founded startups. It finds that women-founded startups had a similar likelihood of securing VC funding within one year of graduation compared to men-founded startups (0.429 vs. 0.443) but raised smaller amounts on average (\$0.352 m vs. \$0.633 m). In terms of performance, women-founded startups had lower failure rates within one year but higher failure rates in the long run.
Table 4 further breaks down fundraising performance by VC funding amount. Note that we observe no difference in the fraction of funded startups across genders (Columns 1 and 2). However, while women-founded startups performed similarly to men-founded startups in percentage at lower funding levels (Column 3), significantly fewer of them raised more than \$2 million both one year and five years from graduation (Columns 4 and 5). [^8] To confirm this, a bootstrapping test in Appendix Table A2 shows no significant difference in mean log-investment size between men- and women-founded startups (p-values: 0.293 at one year, 0.160 at five years). However, the distribution for men-founded startups has significantly higher variance and right skewness, indicating greater access to exceptionally large funding.
Table 3: Startup Performance Post Graduation
All Women Founded Men Founded VC Investments (1 and 5 years) Fraction VC funded (1 year) 0.442 0.429 0.443 Fraction VC funded (5 years) 0.510 0.429 0.517 Investment size (million \$, 1 year) 0.609 0.352 0.633 Investment size (million \$, 5 years) 6.544 1.341 7.031 Operational Status (1 and 5 years) Failure rate (1 year) 0.045 0.032 0.046 Failure rate (5 years) 0.345 0.444 0.336 Acquisition rate (5 years) 0.228 0.206 0.230 Observations 736 63 673 Notes: This table reports average startup outcomes after graduating from an accelerator. Investment size reflects average VC capital raised.
Table 4: VC funding by gender of the founder
(1) (2) (3) (4) (5) (6) Funded Startups Funded Amount Number of Yes No ¿1m ¿2m ¿5m Startups One year after graduation Men founded 375 298 119 46 22 673 (%) (55.7) (44.3) (17.7) (6.8) (3.3) (100.0) Women founded 36 27 9 1 0 63 (%) (57.1) (42.9) (14.3) (1.6) (0.0) (100.0) All 411 325 128 47 22 736 (%) (55.8) (44.2) (17.4) (6.4) (3.0) (100.0) Five years after graduation Men founded 325 348 223 174 123 673 (%) (48.3) (51.7) (33.1) (25.9) (18.3) (100.0) Women founded 36 27 18 12 5 63 (%) (57.1) (42.9) (28.6) (19.0) (7.9) (100.0) All 361 375 241 186 128 736 (%) (49.0) (51.0) (32.7) (25.3) (17.4) (100.0) Notes: This table shows the number and percentage of startups that received VC funding: not funded, funded, funded with more than 1 million US\$, funded with more than 2 million US\$, and funded with more than 5 million US\$.
4 A Matching Model for VCs and Startups
We build a model to discern accelerator and startup quality and to measure the extent of discrimination in this industry. The model features assortative matching, so that high-quality startups and accelerators are more likely to match during admission. We also allow other types of complementarties (e.g., geographic proximity) to affect a match probability. Thus, part of a startup’s unobserved quality can be inferred from the quality of the accelerator it attends.
Importantly for our purposes, our identification strategy allows to separate gender-based sorting across accelerators from gender effects on future startup performances. Suppose a woman-founded startup $s$ and a man-founded startup $s^{\prime}$ are otherwise identical. If gender-based sorting occurs, $s$ and $s^{\prime}$ are unlikely to enter the same accelerator (or those of similar quality) within the same admission market. However, due to market-level variation in competing applicants, they may still end up in the same accelerator. By accounting for other agents in the market, the matching model allows us to compare the unobserved quality of $s$ and $s^{\prime}$. The key identifying assumption is that market composition is exogenous, meaning the unobserved quality of $s$ in a potential match is independent of other agents’ characteristics.
4.1 A Matching Model with Non-transferable Utility
[…]
Agents maximize payoffs by selecting partners on the other side, and equilibrium is defined by pairwise stability, ensuring no agent has a profitable deviation. The equilibrium exists and is unique under stable matching. We estimate the model parameters using a maximum simulated likelihood algorithm.
Market definition.
We define each market as the set of all U.S. accelerators and their program participants within a six-month cohort, beginning in January 2008 and ending in June 2011. The first market includes all matches formed between January and June 2008, the second spans July to December 2008, and so on, through successive six-month periods. This temporal definition avoids imposing strong geographical constraints, unlike prior work—for example, Sørensen, (2007) segments the entire U.S. market into smaller local regions. Given the high incidence of out-of-state matches between startups and accelerators, our approach provides a more flexible and empirically relevant market structure.
We adopt a semiannual market frequency to align with accelerators that run two programs per year, such as Y Combinator. We treat consecutive markets as independent, following Sørensen, (2007) and Fox, (2018), and omit dynamic considerations, as startups typically attend accelerators only once. Sensitivity checks confirm that adjusting market windows does not alter the results qualitatively.
The information structure.
We assume that all startups seeking accelerator programs are aware of all available programs within the same market. This is a standard assumption in the literature, justified by the widespread availability of public announcements online months before admissions begin. However, we do not assume that accelerators have complete knowledge of all potential startups in the market.
Complementarities in portfolio selection.
To ensure the existence of a stable match, we abstract from complementarities in accelerator portfolio selection, assuming that an accelerator’s preference for one startup is independent of its preference for others. Also this assumption is standard in empirical two-sided matching models (e.g., Mindruta et al., , 2016; Akkus et al., , 2020; Pan, , 2017; Honoré and Ganco, , 2020; Akkus et al., , 2016; Fox, , 2018). In practice, accelerators face competition, making deliberate portfolio construction difficult, and it is not uncommon for direct competitors to be admitted to the same cohort (Stross, , 2012).
4.1.1 Setup of the Theoretical Model
Our model has two stages. The first-stage relates to the matching problem between accelerators and startups. The second stage involves startup performances post graduation.
The novelty of our approach lies in keeping the two estimation steps separate. While joint estimation is statistically more efficient, it becomes computationally infeasible in large markets, as the joint likelihood lacks a closed-form expression. As a result, researchers must simulate the distribution over all possible matches, typically restricting analysis to small, local markets (Sørensen, , 2007). However, this limitation may preclude reallocation across distant markets—a key mechanism in our analysis of the gender gap and a quantitatively important channel, with 27% of startups relocating to a different state. By contrast, our two-step approach yields an analytical expression for the matching likelihood in the first step, enabling estimation in settings with many potential matches. In the second step, it allows us to account for match quality using a simple control function approach to study startup performance.
First stage (matching).
[…]
$U_{as}^{s}$ $=(1-E)\times U_{as}$
$U_{as}^{a}$ $=E\times U_{as}$
where $E$ is the (exogenous) equity share of $a$ and is not match-specific. Given $E$, an accelerator $a$ strictly prefers startup $s$ over startup $s^{\prime}$ whenever $U_{as}>U_{as^{\prime}}$, and startup $s$ strictly prefers $a$ over $a^{\prime}$ if and only if $U_{as}>U_{a^{\prime}s}$.
[…]
$\begin{cases}U_{as}>U_{\mu(s),s}\\ U_{as}>\min_{s^{\prime}\in\mu^{-1}(a)}U_{as^{\prime}}.\end{cases}$
That is, $s$ prefers $a$ to its current match $\mu$, and $a$ prefers $s$ to at least one of its current matches in $\mu^{-1}(a)$. A matching $\mu$ is “stable’ if there is no blocking pair. In this context, we can demonstrate that:
Proposition 1.
The accelerator-startup matching model defined under the preferences in 4.1.1 admits a unique equilibrium.
The proof of Proposition 1 follows from applying Proposition 1 in Sørensen, (2007) to our framework. Establishing the existence of a unique stable matching $\mu$ is crucial: without it, the statistical likelihood function is ill-defined, and the absence of a clear equilibrium condition renders the empirical model intractable (Bresnahan and Reiss, , 1991).
Let the observed covariates of $a$ and $s$ be $\mathbf{X}^{as}$. Given the distribution of $\epsilon^{as}$, our first-stage estimator recovers the matching parameters $\beta$ in the expression
$U_{as}=\mathbf{X}^{as}\bm{\beta}+\epsilon^{as}.$ (1)
[…]
Second stage (longer-term performance).
In the second stage, post-matching performance, $Y^{as}$, is modeled as
$Y^{as}=\mathbf{X}^{as}\bm{\alpha}+\eta^{as}.$ (2)
Here, $Y$ can be any startup’s performance after it finishes the accelerator program. In particular, the coefficient of gender is our parameter of interest, and it lies in the coefficient vector $\bm{\alpha}$.
Note that the second-stage error term is $\eta^{as}$ correlated with $\epsilon^{as}$. This is because the VC’s decision $Y$ is also correlated with the unobserved match quality $\epsilon^{as}$, which is observed by the accelerator managers and the startups during the admission process. For each potential pair $(a,s)$, we model the random vector $(\epsilon^{as},\eta^{as})$ using a bivariate normal distribution. Without loss of generality in our context, we normalize the variances so that
$\begin{bmatrix}\epsilon^{as}\\ \eta^{as}\end{bmatrix}\sim\mathcal{N}\left(0,\begin{bmatrix}1&\rho\sigma\\ \rho\sigma&\sigma^{2}\end{bmatrix}\right).$ (3)
This specification captures the idea that higher-quality matches—those with larger values of $\epsilon^{as}$—are more likely to produce stronger post-matching outcomes, reflected in higher $\eta^{as}$. By allowing the two error terms to be correlated, the model accounts for the endogeneity that arises when match quality, observed by the matching agents but unobserved by the econometrician, affects both selection and outcomes. The normalization of variances simplifies identification while preserving the structure needed to estimate the correlation parameter $\rho$.
4.2 Model Estimation
We first estimate the first stage of the matching model and impute the unobservables $\epsilon$ using the estimated parameters. We then control for these unobservables in the second-stage performance equations, enabling the analysis of various outcome variables. This approach also allows us to examine matching patterns independently of second-stage outcomes, including also those startups with no second-stage outcomes (i.e., failed ones).
4.2.1 First Stage: Estimating the Matching Model
[…]
$\epsilon^{as^{\prime}}>\left(\min_{s\in\mu^{-1}(a)}U_{as}\right)-\bar{U}_{as^{\prime}}\ \ \ \ \ \text{ and }\ \ \ \ \ \epsilon^{as^{\prime}}>U_{\mu(s^{\prime}),s^{\prime}}-\bar{U}_{as^{\prime}}$
do NOT hold simultaneously. Define the threshold for blocking as
$\underline{U}_{as^{\prime}}=\max\left\{U_{\mu(s^{\prime}),s^{\prime}}-\bar{U}_{as^{\prime}},\ \left(\min_{s\in\mu^{-1}(a)}U_{as}\right)-\bar{U}_{as^{\prime}}\right\}.$
[…]
$\prod_{a\neq\mu(s^{\prime})}\Phi\left(\underline{U}_{as^{\prime}}\right)$
where $\Phi$ is the c.d.f. of a standard normal distribution due to the assumed marginal distribution of $\epsilon^{as}$ in equation (3). Since this product readily integrates out all unmatched pairs $(a,s^{\prime})$, the overall likelihood of an observed matching $\mu$ is therefore
$\Pr(\mu|X)=\int\left(\prod_{a=\mu(s)}\phi(\epsilon^{as})\right)\left(\prod_{a\neq\mu(s^{\prime})}\Phi\left(\underline{U}_{as^{\prime}}\right)\right)\prod_{a=\mu(s)}\mathrm{d}\epsilon^{as}.$
[…]
4.2.2 Second Stage: Estimating the Performance of Startup Graduates
In the second-stage analysis, we study ex post outcomes $Y^{as}$ through (2). However, $\eta^{as}$ is not independent of $\bm{X}^{as}$ due to its correlation with $\epsilon^{as}$ as displayed in (3). For example, when $a$ and $s$ are distant from each other, they can still form a match if the realized $\epsilon^{as}$ is large enough. Therefore, among the realized matches, distance correlates with $\epsilon^{as}$. As a result, since $\epsilon^{as}$ correlates with $\eta^{as}$, and $\eta^{as}$ correlates with $\bm{X}^{as}$.
[…]
In other words, suppose that we observed $\widehat{\mathbb{E}[\eta^{as}|\mu,\bm{X}]}$, we could use it in (2) to correct for endogeneity problems through a control function approach:
$Y^{as}=\bm{X}^{as}\bm{\alpha}+\widehat{\mathbb{E}[\eta^{as}|\mu,\bm{X}]}+\delta$ (4)
where $\delta:=\eta^{as}-\widehat{\mathbb{E}[\eta^{as}|\mu,\bm{X}]}$. Since the error term $\delta$ now has expectation zero given $\bm{X}^{as}$ (i.e., $\bm{X}^{as}$ is controlled for in $\bm{X}$), this regression can be estimated with no bias through the ordinary least squares (OLS) method, according to:
Proposition 2.
$\mathbb{E}[\eta^{as}|\mu,\bm{X}]$ is a scalar multiple of $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$, i.e., they are collinear.
Proof.
See Appendix section A.4. ∎
The proposition finds that we can use the observed $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$ in place of the unobserved $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$ in (4) to obtain an unbiased least squares estimate for $\bm{\alpha}$. We estimate $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$ by taking the average of the simulated conditional distribution of $\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]$ from the matching model. The standard errors for $\bm{\alpha}$ is obtained through bootstrapping.
4.3 Observables
We hand collect several variables relating to startups, accelerators, and their cohorts.
Gender gap.
To examine the gender gap, we define Female Founder as an indicator for whether at least one among a startup’s founders is female. While analyzing all-women founding teams is of interest, the limited number of such cases constrains our analysis. This measure aligns with the divergence trend observed in Pitchbook data, where gender disparities are likely more pronounced for all-women teams (Box and Segerlind, , 2018).
Team member characteristics.
We collect other characteristics of the founding members with information collected from Linkedin and their personal websites. To capture prior entrepreneurship experience, we define No Serial Founder as an indicator equal to one if no founder has prior startup experience. General work experience, proxied by age, serves as a quality signal, which we examine through Average Age of Founding Team. We also include indicators for educational background: At least One Graduate Degree, At least One PhD Degree, and At least One Engr/Sci Degree. Additionally, Founding Team Size captures the number of founders. Industry differences are accounted for using six-digit 2012 NAICS codes.
Startup characteristics.
Longer-established startups differ from newly founded ones, as they are more likely to have a defined business model, customer base, and experienced founders, all of which facilitate investor attraction. To account for this, we include Startup Age, defined as the number of years since founding prior to accelerator admission.
Accelerator characteristics.
We capture accelerator quality variation with log(Cohort Size), while Accelerator Experience measures the years an accelerator has operated before the current cohort. Accelerator in Startup Hubs indicates whether the accelerator is in MA, CA, or NY. [^10] Data limitations prevent controlling for detailed location fixed effects. To account for potential gender-related differences, ‘All-Men’ Accelerator equals 1 if the founders of the accelerator are all male.
Second-stage performance variables.
We measure startup performance using total VC investment received within one year and five years after demo day. Following the literature, we also track survival status (failed or still operating) and acquisition status (acquired or not) by the fifth year (Hockberg et al., , 2007; Gompers et al., , 2010; Ewens and Rhodes-Kropf, , 2015). Since very few accelerator graduates have completed an IPO, we do not consider IPOs as a measure of performance.
5 Results
We estimate a two-stage model to examine how startup and accelerator characteristics influence both the matching process and subsequent performance outcomes. The first stage models the formation of matches between accelerators and startups. The second stage analyzes post-match outcomes, correcting for endogeneity that arises due to unobserved match quality.
5.1 First Stage: Matching Model
We specify the following linear utility function:
$U_{as}$ $=\beta_{1}\text{Female Founder}_{s}+\beta_{2}\text{``All Men'' Accelerator}_{a}+\beta_{3}\text{Startup Relocated}_{as}$ (5)
$\quad+\beta_{4}(\text{Female Founder}_{s}\times\text{``All Men'' Accelerator}_{a})+\bm{Z}^{as}\bm{\beta}_{Z}+\epsilon^{as},$
[…]
We estimate the parameters in equation (5) via maximum likelihood explained in Section 4.2.1. The computational details are in Appendix A.5. After obtaining parameter estimates $\hat{\bm{\beta}}$, we compute the control function term $\widehat{\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]}$, which captures the conditional expectation of the unobserved quality of a given match, given the realized matching $\mu$ and the vector of observables $\bm{X}$, which spans $\bm{X}^{as}$ for all agents in the market. [^11] Note that $\bm{X}^{as}\subset\bm{Z}^{as}$ because $\bm{X}^{as}$ includes all right-hand side variables in (5).
Appendix A.6 reports the estimation result. The model fits the data well, explaining 78% of the observed and unobserved variation in match qualities. In this section, instead, we focus on the second-stage estimates and the gender gap.
5.2 Second Stage: Outcome Equation
In the second stage, we estimate the effect of startup and accelerator characteristics on a post-match outcome $Y^{as}$, such as whether a startup secures VC funding beyond a given threshold. The estimation equation is:
$Y^{as}=$ $\alpha_{1}\text{Female Founder}_{s}+\alpha_{2}\text{``All Men'' Accelerator}_{a}+\alpha_{3}\text{Startup Relocated}_{as}$ (6)
$+\alpha_{4}(\text{Female Founder}_{s}\times\text{``All Men'' Accelerator}_{a})+\bm{Z}^{as}\bm{\alpha}_{Z}$
$+\lambda\,\widehat{\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]}+\gamma_{t}+\theta_{k}+\delta^{as},$
where $\gamma_{t}$ are year fixed effects capturing year-specific funding environments, $\theta_{k}$ are industry fixed effects based on the startup’s sector, and $\delta^{as}$ is an idiosyncratic error term. The control function term $\widehat{\mathbb{E}[\epsilon^{as}|\mu,\bm{X}]}$ is estimated as the residual of (5) and corrects for endogeneity in the matching process, as justified by Proposition 2.
Fixed effects.
In the second stage, we include year fixed effects $\gamma_{t}$ to account for time-varying funding conditions. These are excluded from the first stage because each market is defined over a six-month cohort, which is more granular than a calendar year, and thus yearly fixed effects cannot be estimated in the first stage.
Industry fixed effects $\theta_{k}$ are included to address concerns that post-match outcomes may differ across tech sub-sectors with varying investment timelines. We do not include these fixed effects in the first stage because accelerators in our sample recruit startups across a wide range of tech industries and generally do not specialize within narrow sub-sectors. Moreover, since accelerators often hold equity stakes for several funding rounds before exit, industry-specific investment cycles are more relevant to performance outcomes than to the initial match formation. Thus, including industry fixed effects in the outcome stage helps absorb any remaining heterogeneity in investment trajectories.
5.3 Post-Accelerator Performance
The first four columns of Table 5 present selected coefficients from estimating (6) using startups’ VC fundraising performance one year after the demo day as dependent variables.
We begin with the one-year horizon because such funding is typically initiated during or shortly after the demo day, as the process from signing the initial contract to receiving the investment often takes several months. [^12] https://www.forbes.com/sites/alejandrocremades/2019/01/03/how-long-it-takes-to-raise-capital-for-a-startup/?sh=1129c4a37a41 Within this short period, the quality of startups changes minimally following graduation, providing us with sufficient control over the startup qualities that are observable to accelerators and investors but not to researchers.
Column (1) uses “Funded” as the dependent variable, an indicator equal to one if the startup secured any VC funding. Columns (2)-(4) examine progressively larger funding thresholds, using indicators for whether a startup raised more than one million, two million, and five million US\$, respectively.
The results in columns (1) and (2) show that the coefficients on “Female Founder” are not statistically significant, suggesting no significant gender differences in the likelihood of receiving any VC funding or amounts below two million US\$. However, in columns (3) and (4), the coefficients on “Female Founder” become significantly negative, indicating that women-founded startups are significantly less likely to raise larger amounts of funding.
Table 5: Second-stage estimation: Performance at one and five years post graduation
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Funding Within One Year Since Graduation Funding Within Five Years Since Graduation Funded? Funding Exceeds (0/1) Funded? Funding Exceeds (0/1) (0/1) One Million Two Millions Five Millions (0/1) One Million Two Millions Five Millions Female Founder 0.053 -0.049 -0.082 -0.038 -0.029 -0.156 -0.109 -0.160 (0.130) (0.102) (0.029) (0.021) (0.123) (0.124) (0.116) (0.073) ”All Men” Accelerator 0.049 0.051 -0.023 -0.009 0.085 0.055 0.051 0.007 (0.062) (0.052) (0.034) (0.025) (0.064) (0.067) (0.061) (0.053) Female Founder $\times$ ”All Men” Accelerator -0.097 0.068 0.066 0.017 -0.097 0.168 0.068 0.103 (0.149) (0.113) (0.039) (0.024) (0.145) (0.143) (0.133) (0.090) Startup Relocated -0.068 -0.105 -0.043 -0.033 -0.046 -0.084 -0.081 -0.034 (0.052) (0.038) (0.022) (0.016) (0.057) (0.053) (0.045) (0.042) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $N$ 736 736 736 736 736 736 736 736 $R^{2}$ 0.093 0.117 0.055 0.034 0.097 0.087 0.056 0.048 Notes: This table presents linear probability estimates from (4), where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
The remaining four columns of Table 5 extend the horizon to funding in the first five years since graduation and find a similar pattern, with large deals being less likely. In particular, the magnitude of the probability of raising at least US\$ 5 million in the first five years is four times larger (in absolute value) than that in column (4), indicating that very few female-founded startups reach this level of financing.
Across columns, we also examine whether “all-men” accelerators further decrease the chances of female entrepreneurs obtaining funding, as prior studies have shown that female investors do not display bias against female entrepreneurs and can help mitigate the gender gap (Raina, , 2021; Hebert, , 2023). The coefficients of “Female Founder $\times$×All Men’ Accelerator” are generally not significant for the probability of being funded (columns 1 and 5). However, it is surprisingly positive and significantive in column (3) for the probability of receiving funding of more than 2 million US\$. Taken together, this evidence does not support the idea that startups with women founders are more likely to match with accelerators with a mixed (or all-women) gender composition.
Finally, the last row of Table 8 highlights the significant costs associated with relocation. Moving to a new state reduces the probability of reaching the one-million and five-million funding milestones by approximately 3.3 to 10 percentage points. To the extent that startups securing five million in funding are of higher quality, their superiority may help them better absorb these relocation costs. Over the five-year horizon, however, the negative impact of relocation diminishes substantially.
In the next section, we further dissect the gender gap to assess the extent to which it stems from relocation costs.
6 What Drives the Gender Gap in Funding?
Thus far, our evidence indicates that a gender gap exists, with female founders having lower chances of receiving large amounts of funding. In the following section, we relate this result to relocation costs.
A significant difference highlighted in the literature is that women face higher relocation costs than men (Bielby and Bielby, , 1992; Jayachandran et al., , 2024). In particular, women founders may be less likely to move their startups to a different state even if they expect a better match for their startup with an accelerator in that state. If this factor explains the observed funding gap, women-founded startups that relocated to a different state to join accelerators should raise more funding after graduation.
To view the role of relocating challenges, Table 6 interacts “Female Founder” and the indicator “Startup Relocated,” which is 1 if the startup relocated to a new state and zero otherwise. As in the previous table, the first four columns report startups’ performance one year after graduation, while the remaining columns extend the analysis to five years. Importantly, because of our control function approach (Proposition 2) we can consider startup relocation as exogenous as relocation was controlled for in the matching utilities (1)—as for all the other covariates in (4)—ensuring a causal interpretation of our findings.
Across columns, the estimated coefficients for “Female Founder” closely align with those in Table 5. For example, female-founded teams have a 7.6 percentage point lower probability of securing at least two million US\$ in the first year after graduation (Column 3) and a 17.9 percentage point lower probability of raising five million US\$ within five years (Column 8). However, this gap narrows for startups willing to relocate, as the coefficient in the second row is generally positive despite large standard errors due to the small sample. This result suggests that relocation mitigates the gender gap in funding.
To further assess this mechanism, the table reports a two-sided test for whether the sum of the gender gap coefficient and its interaction term with “Startup Relocated” equals zero. Across columns, we fail to reject the null hypothesis in all cases except Column (3), where the dependent variable is the probability of securing two million US\$ within one year of graduation.
Therefore, this null result provides suggestive evidence that of the mechanism underlying the gender gap in Table 5: not all startups may be able to match with their favorite accelerator, leading to suboptimal performance and inefficiencies.
Table 6: The role of relocation
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Funding Within One Year Funding Within Five Years Funded? Funding Exceeds (0/1) Funded? Funding Exceeds (0/1) (0/1) One Mil. Two Mil. Five Mil. (0/1) One Mil. Two Mil. Five Mil. Female Founder ($\alpha_{1}$) -0.014 -0.078 -0.076 -0.041 -0.087 -0.195 -0.111 -0.179 (0.135) (0.101) (0.030) (0.023) (0.132) (0.129) (0.126) (0.086) Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.238 0.102 -0.022 0.010 0.209 0.139 0.007 0.065 (0.143) (0.116) (0.034) (0.019) (0.143) (0.146) (0.129) (0.106) Other Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$ 0.225 0.024 -0.098 -0.031 0.122 -0.056 -0.104 -0.113 (0.149) (0.150) (0.039) (0.029) (0.144) (0.164) (0.140) (0.103) $N$ 736 736 736 736 736 736 736 736 $R^{2}$ 0.097 0.118 0.055 0.034 0.100 0.088 0.056 0.048 Notes: This table presents linear probability estimates from (4) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
Confounding factors.
Higher-quality startups may be more likely to relocate, raising the concern that our analysis captures unobserved aspects of startup quality correlated with female-founded teams rather than the actual match quality and complementarity between startups and accelerators, despite our control function approach.
To address this issue, we replicate the previous analysis using a restricted sample that includes only startups that survived their first year. This approach removes the lowest-quality startups, creating a more homogeneous sample at the cost of a smaller dataset and larger standard errors. Appendix Table A3 presents the results.
Across columns, the point estimates for the gender gap (first row) are even larger than those in Table 6. Meanwhile, the estimated benefits of relocating appear smaller. Although the test never rejects the null hypothesis that the sum of the two coefficients is zero, this result may be driven by the larger standard errors due to the smaller sample. Notably, the point estimate for the sum of “Female Founder” and “Female Founder $\times$ Startup Relocated” is larger in magnitude than in Table 6.
Since, if higher-quality startups were to be more likely to relocate we would have expected larger gains from relocation, this analysis reinforces the presence of a gender gap in funding and suggests that female-founded startups willing or able to relocate can partially close this gap.
6.1 Mechanism
Costs for mother founders. Relocation may be particularly costly for mother entrepreneurs. Relocating with a newborn or a young child may be particularly difficult, regardless of the quality of the startup.
Unfortunately, we do not observe whether a founder is a mother. Moreover, controlling for parenthood alone may be insufficient, as plans to have children can also constrain mobility. While some administrative data might capture parenthood and surveys could reveal fertility intentions, these variables are unavailable in most datasets. Thus, to explore whether the gender gap is more pronounced in cases where family constraints are more likely, we restrict the sample to startups whose teams have an average age between 28 and 40. This range corresponds to typical childbearing ages among highly educated women: the median age at first birth for degree-holding women is around 28, and tends to be even higher for those in STEM fields—relevant for our sample of startup founders (Livingston, , 2015; Lappegard et al., , 2020; Schweizer and Guzzo, , 2020).
We present the results in Table 7, which focuses on the probability of observing fundings exceeding two and five million US\$ at either horizon, as the previous analyses highlight that the gender gap is especially pronounced for these funding levels, despite larger standard errors arising from the smaller sample size.
The results indicate a larger gender gap (row 1) compared to the estimates obtained from the full sample. The gender gap appears particularly large over the five-year horizon (Columns 3 and 4). For instance, female founders who do not relocate are 34 percentage points less likely to reach both the two- (non-significant) and the five-million funding milestones. This result comes as no surprise, as parenting may take significant resources from mother entrepreneurs, and the time they allocate away from their startups has a substantial economic cost five years after graduation.
The second line offers new insights into the benefits of relocation. While no clear advantage emerges over short horizons (columns 1 and 2), relocation is associated with substantial gains over the five-year horizon. Relocating female founders are 26.5 percentage points more likely to exceed the two-million funding threshold and 41 percentage points more likely to surpass five million. The latter effect is significant at the 10% level, and the point estimate exceeds the corresponding gender gap in the first row—suggesting that relocation can potentially offset the gender gap.
Notably, the fact that the largest effect appears at the five-million threshold, and only over five years, suggests that it may take time for female-founded startups in this age bracket to catch up—possibly due to support mechanisms implemented by the accelerator. In the next section, we examine how accelerators could help ease relocation costs.
Table 7: The role of relocation for female entrepreneurs in motherhood age
Linear Probability Model: (1) (2) (3) (4) Funding Within One Year Funding Within Five Years Exceeds (0/1) Exceeds (0/1) Two Millions Five Millions Two Millions Five Millions Female Founder ($\alpha_{1}$) -0.113 -0.077 -0.335 -0.344 (0.057) (0.045) (0.212) (0.117) Female×$\times$ Startup Relocated ($\alpha_{2}$) -0.062 0.015 0.265 0.410 (0.077) (0.045) (0.255) (0.229) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$0.175 -0.062 -0.070 -0.067 (0.085) (0.064) (0.260) (0.229) Sample Average Age in [28-40] years $N$ 337 337 337 337 $R^{2}$ 0.086 0.077 0.109 0.132 Notes: This table presents linear probability estimates from (4) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Only startup whose founders’ average age is between 28 and 40 are considered. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
Can accelerators level the gap? We leverage variation across accelerators to examine the gender gap at higher-quality accelerators.
Table 8 presents regression results on another subsample. The first four columns focus on networking effects: we restrict the sample to accelerators with cohorts larger than the median, under the assumption that larger accelerators—such as Y Combinator—have more resources, can host more startups, and generate stronger network externalities among participants.
Once again, our findings reveal a stark contrast between short- and long-term funding outcomes. The estimated gender gap (row 1) is similar to the full-sample estimates from columns (3) and (4) of Table 6. Moreover, the benefits of relocation appear limited, particularly in column (1). However, columns (3) and (4) offer a different perspective over the five-year horizon: the gender gap narrows substantially, shrinking by approximately 60% in column (3) $\left(\frac{-0.044+0.111}{-0.111}\right)$ and by 39% in column (4) $\left(\frac{-0.110+0.179}{-0.179}\right)$. The gender gap among relocating female founders decreases even further, to about one-fourth and half of the values observed in Table 6.
These results suggest that one key advantage of larger cohorts is that by engaging with a broader group of founders, relocating female entrepreneurs build a support network that fosters their startups’ growth. Further research could explore the underlying mechanisms driving this effect. Additionally, a stronger network may allow them to overcome time constraints (as the coefficients on the first row remain negative) by providing access to trusted individuals who can assist with specific business challenges. Enhanced networking opportunities may also facilitate strategic partnerships between startups or among key individuals within the ecosystem.
The remaining columns of Table 8 examine a sample of more experienced accelerators. While the gender gap persists for one-year funding milestones (columns 5 and 6), it fully disappears over the five-year horizon (columns 7 and 8). As the investment horizons of accelerator is approximately five years, experienced accelerators may offer structured mentorship programs tailored to female founders, equipping them with the skills necessary to successfully scale their startups and catch up by this date (Zhang, , 2023).
Table 8: The role of accelerators at closing the “relocation” gap
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Accelerators with larger cohorts More experienced accelerators Funding Within One Year Funding Within Five Years Funding Within One Year Funding Within Five Years Exceeds (0/1) Exceeds (0/1) Exceeds (0/1) Exceeds (0/1) Two Mil. Five Mil. Two Mil. Five Mil. Two Mil. Five Mil. Two Mil. Five Mil. Female Founder ($\alpha_{1}$) -0.072 -0.044 -0.044 -0.110 -0.093 -0.036 0.000 0.125 (0.043) (0.033) (0.149) (0.103) (0.043) (0.031) (0.303) (0.297) Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.018 0.037 0.069 0.053 -0.063 -0.041 0.262 0.342 (0.055) (0.039) (0.217) (0.162) (0.126) (0.081) (0.429) (0.406) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$0.053 -0.006 -0.025 -0.056 -0.156 -0.076 0.262 0.466 (0.062) (0.050) (0.196) (0.137) (0.141) (0.096) (0.515) (0.493) Sample Large cohorts (above median) Experienced accelerator (above median) $N$ 408 408 408 408 326 326 326 326 $R^{2}$ 0.082 0.061 0.060 0.071 0.076 0.060 0.098 0.088 Notes: This table presents linear probability estimates from (4) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Columns 1 to 4 focus on startups in accelerators with cohorts above the median cohort size. Columns 5 to 8 focus on startups in accelerators with more than the median number of years of activity. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
We acknowledge that the confidence intervals for these estimates are quite large, reflecting the small sample size. [^13] To the extent that the population of interest consists of firms matching with accelerators in the U.S. during our study period, we observe the entire population. In this casee, our standard errors are conservative (Sancibrián, , 2024). This limitation prevents us from further expanding the sample to explore additional mechanisms at play. Nevertheless, our findings underscore the importance of complementarities between startups and accelerators, pointing to promising avenues for future research. In particular, leveraging datasets that track interactions between founders within an accelerator, as well as the specific programs available, could provide deeper insights into how accelerators help reduce the gender gap in VC financing.
6.2 Economic Benefits of Bridging the Gender Gap
In this section, we document that policies aimed at reducing the gender gap in financing are beneficial not only from a fairness perspective but also in economic terms. This is because startups founded by women are not less productive than comparable startups. Although measuring productivity in startups is challenging due to the lack of reliable revenue and employment data, we use the probability of failure in the early years after graduation as a proxy for productivity.
To assess this, we analyze the first three columns of Table 9. Column (1) examines failure within the first year, while column (2) considers failure within the first five years. The results indicate that female-founded startups are not significantly more likely to fail than their counterparts. While these measures capture the “lack of productivity,” column (3) presents a complementary perspective, showing that female-founded startups are equally likely to be acquired or go public (IPO) within five years—an indicator of high productivity, as these outcomes often signal successful and profitable exits.
The remaining columns further investigate these findings by conditioning on surviving startups. Columns (4) to (6) confirm that female-founded startups receive less funding, particularly struggling to reach the five-million milestone. Despite this funding gap, female-founded startups are just as likely to survive or successfully exit within five years as other startups. This suggests that female-founded startups can achieve similar outcomes as male-founded startups with fewer financial resources, highlighting their high productivity.
As a result, further bridging the gender gap in financing could not just merely increase the number of startups but also create more successful ones. While additional research is needed to better understand the specific policies implemented by higher-quality accelerators, our findings suggest that accelerators sustaining diversity among startup founders can grow the pie, being beneficial for both firms and society.
Table 9: Operating status
Linear Probability Model: (1) (2) (3) (4) (5) (6) (7) (8) Failed in its First Acquired Funding Exceeds (0/1) Failed Exited 1 Year 5 Years in 5 Years One Million Two Millions Five Millions (0/1) (0/1) Female Founder 0.068 0.029 0.068 -0.185 -0.133 -0.213 -0.026 -0.015 (0.128) (0.137) (0.134) (0.186) (0.184) (0.129) (0.141) (0.162) Other Controls $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Sample All All All Conditional on surviving in the first year since graduation $N$ 736 736 736 325 325 325 325 325 $R^{2}$ 0.044 0.108 0.044 0.091 0.061 0.091 0.072 0.055 Notes: This table presents linear probability estimates from (4) of startup outcomes, including failure, acquisition, and funding milestones. The first three columns analyze failure within one and five years, as well as acquisition within five years. Columns (4)–(6) examine whether a startup raised more than \$1 million, \$2 million, or \$5 million, conditional on surviving the first year. The last two columns report failure and exit probabilities. Bootstrapped standard errors are in parentheses. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors in parentheses.
7 Conclusions
This study provides insights into the gender gap in venture capital funding, focusing on how gender-based disparities persist in larger funding rounds despite similar early-stage performance. By analyzing data from accelerator graduates, we highlight that female-founded startups face significant challenges, particularly due to the relocation constraints often associated with family responsibilities. Our findings suggest that while accelerators play a crucial role in fostering startup growth, they also have the potential to mitigate the gender gap in funding. Reducing this gap is important not only for fairness but also for maximizing the economic potential of women entrepreneurs.
We also contributed by offering a computationally light approach to estimating matching designs in large markets. In particular, the methodology used in this paper combines a one-to-many matching model with non-transferable utility and a control function approach to estimate startup outcomes. By first modeling the matching process between startups and accelerators, and then using this information to control for unobserved startup quality, we provide a robust analysis of funding disparities. This approach allows us to account for the endogenous factors that may affect venture outcomes and offers a nuanced understanding of the gender funding gap in the venture capital landscape.
Appendix A Online Appendix
A.1 Accelerator Process
As shown in Figure A1, the accelerator procedure starts with a public announcement of the details and terms of the program, including information such as cohort size, location, and schedule. Once announced, these terms rarely change and are not subject to negotiation. Startups submit their applications to the accelerators that they would like to join, and the accelerators admit the strongest applicants based on predetermined cohort capacities. Admitted entrepreneurs start the program together at the same time and in the same location. The program lasts for a fixed period, often three months, during which the accelerator offers mentorship, network opportunities, and other business support. At the end of the program, the accelerator invites potential investors to join a “demo day” during which the graduating startups present their pitches. The graduating startups pitch to investors to secure funding. The participating firms are under no obligation to the accelerator after graduation, but they often remain involved in the community as alumni.
Figure A1: Accelerator Process
A.2 Gender Gaps in the VC Market
Figure A2 shows the decreasing gender gap in terms of the number of VC deals. The curve shows a downward trend in the difference between the number of VC deals obtained by startups with all-men founders and the number of VC deals obtained by startups with all-women founders.
Figure A2: Gender gap in the relative number of VC deals
[…]
Figure A3 shows the increasing gender gap in terms of average funding size. In the figure, women-founded startups means there is at least one woman on the founding team; all-women startups is defined as startups with only female founders; women-led startups is defined as startups whose CEO is a woman; and women-founded startups in Tech is defined as startups in tech-industry whose founders include at least one woman.
Figure A3: Diverging gender gap in average investment sizes of VC deals
Note: This figure shows the differences in the average VC investment sizes between all-men founded startups and 1) women-founded startups, 2) all-women startups, and 3) women-led startups (female CEO). In addition, it also shows the gap in average investment sizes between all-men and women-founded startups in Tech. The Y axis is in unit of million US\$.
A.3 Omitted Tables
Table A1: Accelerator admission and political environment
Indicator for Share of Female Founded Startups Female Founded Sturtups (1) (2) (3) (4) Acc in Democratic State 0.007 0.012 0.025 0.028 (0.032) (0.032) (0.032) (0.033) Year Fixed Effects Y Y $N$ 736 736 74 74 $R^{2}$ 0.000 0.002 0.009 0.022 Notes: This table shows the coefficients from regressing the occurrence of female founded startups on a dummy for voting for the democratic party in the 2008 presidential election. The dependent variables for the first two models are indicators of whether the accelerators’ participating startups are founded by women. The dependent variables for the last two models are the percentages of women-founded startups. Even columns include state-fixed effects.
Table A2: Startup VC funding
Women Men p-value Founded Founded of W¿M log(InvestSize 1yr) Mean 6.427 6.562 0.293 Variance 0.777 1.624 0.027 Skewness -1.065 -0.397 0.114 log(InvestSize 5yr) Mean 7.388 7.768 0.160 Variance 1.602 3.768 0.004 Skewness -0.204 0.048 0.032 Notes: This table shows the log(InvestSize) distribution moments for VC funded startups. The last column presents the p value of whether the corresponding moment for women-founded startups is larger than that of their male counterparts.
Table A3: The role of relocation for surviving startups
Linear Probability Model: (1) (2) (3) (4) (5) (6) Funding Within One Year Funding Within Five Years Exceeds (0/1) Exceeds (0/1) One Mil. Two Mil. Five Mil. One Mil. Two Mil. Five Mil. Female Founder ($\alpha_{1}$) -0.125 -0.130 -0.065 -0.174 -0.044 -0.206 (0.182) (0.076) (0.053) (0.202) (0.198) (0.165) Female×$\times$ Startup Relocated ($\alpha_{2}$) 0.113 -0.022 0.041 -0.026 -0.214 -0.018 (0.202) (0.086) (0.048) (0.206) (0.230) (0.188) Other Control Variables $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Correction Term $\mathbb{E}[\eta^{as}|\mu,X]$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Year Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Industry Fixed Effects $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Test $H_{0}:\alpha_{1}+\alpha_{2}=0$0.012 -0.152 -0.024 -0.199 -0.257 -0.223 (0.220) (0.088) (0.066) (0.215) (0.207) (0.159) Sample Conditional on surviving in the first year since graduation $N$ 325 325 325 325 325 325 $R^{2}$ 0.161 0.103 0.062 0.091 0.064 0.091 Notes: This table presents linear probability estimates from (4) adding the interaction term between Female×$\times$ Startup Relocated, where the dependent variables measure performance within one year and cumulatively by year five after graduation. They indicate whether a startup received VC funding, more than \$1 million, more than \$2 million, or more than \$5 million. The Test panel tests the null hypothesis that the sum of the coefficients for Female Founder ($\alpha_{1}$) and Female×$\times$ Startup Relocated ($\alpha_{2}$) is zero. Only startup that survived the first year after graduation are considered. All specifications include controls (see Section 4.3 for details), correction terms, and fixed effects as indicated. Bootstrapped standard errors are in parentheses.
A.4 Proof of proposition 2.
Proof.
Observe that through law of iterated expectation, we have
$\mathbb{E}[\eta^{as}|\mu,X]=$ $\mathbb{E}\left[\mathbb{E}[\eta^{as}|\epsilon,\mu,X]|\mu,X\right]$
$=$ $\mathbb{E}\left[\mathbb{E}[\eta^{as}|\epsilon,X]|\mu,X\right]$
$=$ $\mathbb{E}\left[\mathbb{E}[\eta^{as}|\epsilon^{as}]|\mu,X\right]$
$=$ $\mathbb{E}\left[\rho\sigma\epsilon^{as}|\mu,X\right]$
$=$ $\rho\sigma\mathbb{E}\left[\epsilon^{as}|\mu,X\right]$
The second equality is due to the fact that the $\sigma$-field generated by $(\epsilon,X)$ determines $\mu$. The third equality is due to the fact that $(\epsilon^{a^{\prime}s^{\prime}},X)$ are independent of $\eta^{as}$ when $a^{\prime}s^{\prime}\neq as$. The fourth equality follows from properties of bivariate normal distribution. Consider the product $\rho\sigma$ as a deterministic parameter different from zero; this completes the proof. ∎
A.5 Simulated Maximum Likelihood Algorithm
The steps taken to perform the maximum simulated likelihood estimation are detailed below. Suppose that there are $K$ markets $\{1,\dots,K\}$ where the $k$th market has observed matching $m_{k}$ that contains $|m_{k}|$ number of matched pairs.
  1. 1.
    For the $k$th market with $|m_{k}|$ matched pairs $\{(a,s)_{i}\}_{i=1}^{|m_{k}|}$, simulate vectors $\epsilon^{as}$ from an i.i.d normal distribution of dimension $m_{k}$. Independently simulate a large number $T$ of such $\epsilon$-vectors, e.g., $T=10000$.
  2. 2.
    For the $k$th market, with observed matching $m_{k}$, $g_{k}(\beta,\epsilon^{as})=\sum_{a^{\prime}\neq m_{k}(s^{\prime})}\ln\Phi(\underline{U}_{a^{\prime}s^{\prime}})$. where $\underline{U}_{a^{\prime}s^{\prime}}$ is defined as in the main text. Here, $g_{k}$ is a function of the parameters of interest and the $|m_{k}|$-dimensional vector $\epsilon^{as}$.
  3. 3.
    Choose $\beta$ to maximize the objective
    $LogSumExp\left(g(\beta,\epsilon^{as}_{t})\right)=\ln\left(\sum_{k=1}^{K}\sum_{t=1}^{T}\exp\left[g_{k}(\beta,\epsilon^{as}_{t})\right]\right).$
    The solution is our point estimate $\hat{\beta}$.
A.6 The Matching Model Estimates
Table A4 reports the estimates of the matching model. Column Coef is the $\beta$ as in the match value function of Equation 1. We also report the standard errors, obtained from bootstrapping, of our point estimates. In addition to all of the empirical controls discussed in Section 4.3, we include an additional indicator Startup Relocated to capture whether the startup had to relocate to a different state to join the accelerator. Such relocation can be very costly for a startup, not only because the founding team needs to change its place of residence but also because the startup might lose its original local support, business partner(s), and customer base.
Table A4: First-stage results: Admission matching
Coefficient Std Err
Female Founder -0.326 0.142
No Serial Founder -0.124 0.215
Startup Age -0.198 0.303
At least One Graduate Degree 0.119 0.210
At least One PhD Degree 0.196 0.163
At least One Engr/Sci Degree 0.700 0.230
Average Age of Founding Team 0.020 0.042
Founding Team Size -0.353 0.235
Accelerator in Startup Hubs (CA, NY, MA) -0.261 0.162
Accelerator Experiences (yrs) 0.060 0.040
log(Cohort Size) 0.652 0.119
Accelerator w Female Founder 0.132 0.160
Accelerator w Female Founder×$\times$×Female Founder -0.006 0.165
Startup Relocated -2.692 0.113
Our matching model estimates indicate that startups founded by women are less valued in the accelerator market, as indicated by the negative parameter for Female Founder. To measure the goodness-of-fit for the first-stage matching model, we compare the variance of $X^{as}\hat{\beta}$ from the structural component of the matching value, to the variance of $\hat{\epsilon}^{as}$ from the imputed unobserved matching quality. Because the value of a match is determined according to the model as $U^{as}=X^{as}\beta+\epsilon$, this comparison provides a measure analogous to the multiple $R^{2}$ in a regression. We find that $Var[X^{as}\hat{\beta}]/Var[\hat{\epsilon}]=4.63$, comparable to an $R^{2}$ of approximately 82%.